# One Shot Dominance: Knowledge Poisoning Attack on Retrieval-Augmented Generation Systems

### v1

@@ 4.4 Baselines
  • PoisonedRAG (PRAG) Zou et al. (2024): Uses LLM to generate supporting content prefixed with complete questions to ensure high retrieval probability.
@@ 4.5 Experiment Design and Metric
For the approach implementation, we leverage GPT-4 as the backbone for the Content Agent, the CoE Judge Agent, the revise agent, and the Authority Content Agent, consistent with the original PoisonedRAG work where GPT-4 was used to generate poisoned data Zou et al. (2024). The temperature parameter is set to 0.1 for all agents to ensure the stability and consistency of generated content.
To answer RQ1, we inject a single poisoned document, generated by AuthChain, into the knowledge base for each targeted question and evaluate its effectiveness in manipulating the outputs of various LLMs under a RAG setting. We compare AuthChain against baselines, while also evaluating its internal components (CoE and Authority content) to provide a comprehensive performance analysis. The evaluation is conducted using three metrics: Attack Success Rate (ASR), Retrieval Success Rate RSR, and Perplexity (PPL). To measure the ASR, we calculate the proportion of questions for which the LLM’s output contains the answer from the poisoned target document among all questions. Following previous works Rizqullah et al. (2023); Huang et al. (2023), we determine whether the target answer is present in the LLM’s output using a substring matching approach. For RSR, we compute the proportion of poisoned target documents that are successfully retrieved among the top-5 documents for all poisoning targets. As for PPL we utilize GPT-2 Radford et al. (2019) to calculate the text perplexity score, which measures text fluency. Higher PPL values indicate less natural and more anomalous text generation.
@@
To answer RQ2, we evaluate how effectively the poisoned documents generated by AuthChain can evade detection mechanisms within RAG-based defense frameworks. Specifically, we select two representative RAG defense frameworks designed to counteract knowledge poisoning attacks: InstructRAG Wei et al. (2024) and AstuteRAG Wang et al. (2024b). InstructRAG enhances the robustness of RAG systems by explicitly guiding language models to learn a denoising process based on self-synthesized rationales. In this framework, the model is instructed to explain how the ground-truth answer is derived from the retrieved documents. These rationales can be leveraged as in-context demonstrations for explicit denoising or as supervised fine-tuning data, thereby improving the model’s ability to identify and resist poisoned or misleading knowledge in the retrieval set. AstuteRAG improves the robustness of RAG systems against imperfect or malicious retrieval by analyzing conflicts between the LLM’s internal knowledge and external sources. It adaptively extracts key information from internal knowledge, integrates it with retrieved content, and produces answers based on source reliability. This method has shown strong effectiveness in detecting and mitigating knowledge poisoning attacks. We evaluate the effectiveness of these frameworks using two metrics: Attack Success Rate (ASR) and Accuracy (ACC). ASR is the same as defined in RQ1. ACC is defined as the proportion of questions for which the LLM’s response contains the correct answer, among all questions. We evaluate both AuthChain and baselines under these two defense frameworks, comparing with the clean setting where no poisoning attack is applied.
To answer RQ3, we conduct two experiments: (1) Authority Setting: We randomly sample 600 QA pairs from HotpotQA and identify 118 questions where GPT-3.5 demonstrates clear internal knowledge by correctly answering without external retrieval. As GPT-3.5 was one of the earliest GPT models, later versions in the GPT series not only retain this internal knowledge but also correctly answer these 118 questions, making them ideal candidates for our verification. For these questions, we retrieve supporting documents from the knowledge base and manually alter their answers to incorrect ones, creating initial poisoned documents (defined as Raw documents). We further enhance these raw documents by automatically inserting authoritative statements generated by AuthChain. We then gradually introduce correct supporting documents to construct external knowledge with correct document proportions (CDP) of 0.5, 0.67, and 0.75, and measure the effectiveness using ASR, as defined in RQ1. We evaluate the impact of authority statements by comparing the ASR between raw documents and authority enhanced documents. (2) CoE Setting: Working on the same 600 QA pairs, we identify 323 support documents provided in the dataset that contain evidence for correct answers but lack CoE features (defined as Raw documents). We convert these Raw documents into CoE content using AuthChain. We then use the GPT-4 to generate poisoned documents that contain evidence leading to incorrect answers for these questions. These poisoned documents are incrementally introduced to construct external knowledge with poisoned document proportions (PDP) of 0.5, 0.67, and 0.75. The impact is measured using ACC, the same as defined in RQ1. We evaluate the effectiveness of CoE by comparing the ACC between raw documents and CoE structured documents.
@@ 5.1 Answering RQ1
We evaluate the effectiveness of AuthChain against six popular LLMs (GPT-3.5, GPT-4, GPT-4o, Llama3-8B, Llama3-70B, and DeepSeek-V3) across three widely-used datasets (HotpotQA, MS-MARCO, and NQ). Table 1 compares the Attack Success Rate (ASR) of our method with state-of-the-art baselines (PRAG and HRAG), along with key metrics including Retrieval Success Rate (RSR) and perplexity (PPL).
AuthChain achieves an average ASR of 87.0%, 81.5%, and 77.8% on HotpotQA, MS MARCO, and NQ respectively, surpassing PRAG (21.7%-36.7%) and HRAG (31.0%-46.5%) by a significant margin. This demonstrates the effectiveness of AuthChain where a single poisoned document can effectively dominate other retrieved knowledge and gain LLM preference. As for RSR, PRAG and HRAG concatenate the entire question into poisoned document to increase retrieval probability, while AuthChain guides LLMs to generate poisoned content with more CoE features, resulting in more natural and fluent text. This improved fluency is quantitatively validated by the PPL, where AuthChain achieves significantly lower PPL (average 33.4) compared to PRAG (69.9) and HRAG (381.1). Notably, AuthChain maintains competitive retrieval performance with only an average 2.3% decrease in RSR compared to PRAG and an average 5.0% higher RSR than HRAG, while achieving substantially higher ASR.
Both components contribute significantly to this improvement. The CoE component achieves an average 10.2%-18.8% higher ASR than baselines, as its structured knowledge better aligns with questions and proves more effective at influencing LLM’s external knowledge selection. The Authority component shows an average 10.0%-18.6% improvement over baselines, revealing the widespread authority bias vulnerability inherent in LLMs. This susceptibility persists across different LLM architectures and scales, indicating a fundamental weakness in current LLM systems rather than an isolated phenomenon. When combined, AuthChain surpasses both individual components by 18.4% and 18.6% respectively, suggesting that CoE and Authority complement each other by targeting different aspects: CoE enhances external knowledge comprehension while Authority challenges internal knowledge robustness.
Notably, GPT-4o demonstrates stronger defense capabilities, as shown by its significantly lower ASR against baseline attacks compared to other tested LLMs, reflecting its enhanced ability to identify and resist misleading knowledge. Similarly, Llama3-70B exhibits remarkable resilience specifically against HRAG’s prompt injection attacks, with extremely low ASR (19.3%), suggesting this LLM has been specifically fine-tuned to defend against prompt injection attacks. Nevertheless, AuthChain maintains consistent ASR across all tested LLMs by generating content that more naturally resembles authentic text. To alleviate the computational costs of using GPT-4 based agents, we explore open-sourced LLMs as alternative agents in Appendix G. The results show these alternatives achieve comparable performance while being more cost-effective.
@@
Dataset Model (ASR) PRAG HRAG CoE Authority AuthChain
HotpotQA GPT-3.5 69.0 57.0 79.0 82.0 90.0
@@
RSR 97.0 87.0 93.0 65.0 94.0
PPL 73.8 398.0 28.5 49.5 26.4
@@ 5.2 Answering RQ2
Defense Method Dataset Model PRAG HRAG AuthChain Clean
[…]
InstructRAG HotpotQA GPT-3.5 49.0/42.0 40.0/46.0 36.0/60.0 76.0
GPT-4 56.0/38.0 52.0/39.0 47.0/52.0 79.0
GPT-4o 68.0/31.0 70.0/24.0 58.0/40.0 81.0
Llama3-8B 53.0/42.0 48.0/44.0 47.0/50.0 82.0
Llama3-70B 66.0/30.0 78.0/14.0 60.0/38.0 84.0
Deepseek-V3-0324 69.0/30.0 62.0/32.0 60.0/36.0 80.0
MS-MARCO GPT-3.5 51.0/40.0 47.0/43.0 45.0/46.0 78.0
GPT-4 65.0/30.0 59.0/39.0 56.0/43.0 86.0
GPT-4o 76.0/21.0 81.0/13.0 70.0/25.0 82.0
Llama3-8B 56.0/37.0 49.0/48.0 51.0/46.0 83.0
Llama3-70B 67.0/29.0 71.0/22.0 50.0/48.0 79.0
Deepseek-V3-0324 82.0/16.0 72.0/20.0 69.0/28.0 89.0
NQ GPT-3.5 49.0/42.0 45.0/50.0 41.0/56.0 66.0
GPT-4 64.0/29.0 54.0/40.0 51.0/44.0 75.0
GPT-4o 83.0/14.0 75.0/18.0 69.0/27.0 78.0
Llama3-8B 54.0/38.0 50.0/40.0 49.0/44.0 74.0
Llama3-70B 65.0/30.0 72.0/16.0 47.0/45.0 79.0
Deepseek-V3-0324 79.0/21.0 69.0/28.0 68.0/30.0 82.0
AstuteRAG HotpotQA GPT-3.5 59.0/33.0 61.0/25.0 46.0/52.0 78.0
GPT-4 79.0/10.0 74.0/13.0 63.0/28.0 78.0
GPT-4o 72.0/11.0 78.0/10.0 52.0/40.0 79.0
Llama3-8B 70.0/15.0 52.0/31.0 46.0/48.0 70.0
Llama3-70B 72.0/17.0 76.0/8.0 51.0/45.0 83.0
Deepseek-V3-0324 78.0/8.0 71.0/15.0 60.0/34.0 75.0
MS-MARCO GPT-3.5 73.0/14.0 81.0/12.0 56.0/30.0 82.0
GPT-4 88.0/5.0 85.0/8.0 57.0/34.0 89.0
GPT-4o 86.0/4.0 84.0/8.0 67.0/28.0 87.0
Llama3-8B 83.0/11.0 73.0/23.0 60.0/35.0 91.0
Llama3-70B 86.0/9.0 58.0/39.0 56.0/37.0 91.0
Deepseek-V3-0324 88.0/5.0 85.0/10.0 82.0/11.0 89.0
NQ GPT-3.5 62.0/21.0 75.0/9.0 44.0/47.0 71.0
GPT-4 81.0/5.0 83.0/7.0 69.0/23.0 80.0
GPT-4o 80.0/8.0 82.0/9.0 71.0/23.0 84.0
Llama3-8B 78.0/9.0 69.0/17.0 60.0/33.0 78.0
Llama3-70B 83.0/5.0 87.0/5.0 58.0/38.0 86.0
Deepseek-V3-0324 86.0/0.0 84.0/6.0 75.0/11.0 86.0
@@ 5.3 Answering RQ3
Table 3 presents results from two distinct experiments examining the effectiveness of Authority and CoE components. The first experiment investigates the Attack Success Rate (ASR) with and without authority statements as the Correct Document Proportion (CDP) increases. The second experiment evaluates the impact of CoE on LLM accuracy (ACC) under varying Poisoned Document Proportion (PDP).
For the authority setting, with raw poisoned documents (without authority statements) and CDP set to 0.5, LLMs show considerable internal robustness, resulting in relatively low ASR (average 32.2%). This is particularly evident in advanced models like GPT-4o, where the ASR is only 17.9%, demonstrating strong internal knowledge resistance to poisoning. As CDP increases, the ASR of raw poisoned documents drops even further. However, incorporating authority statements significantly improves attack effectiveness, achieving an average ASR of 60.9% with CDP of 0.5, marking a 28.7% increase over raw poisoned documents. Moreover, as CDP rises to 0.75, documents with authority statements maintain better effectiveness, showing only a 15.6% ASR decrease compared to the 23.3% decrease in raw documents. This demonstrates that authority statements effectively overcome LLMs’ internal knowledge barriers and enhance the credibility of poisoned content.
Regarding the CoE experiment, supporting documents structured with CoE help LLMs achieve an average ACC of 88.1% compared to 79.6% with raw supporting documents when PDP is 0.5. Even as PDP rises to 0.75, CoE supporting documents maintain an ACC of 83.2%, while raw documents drop to 72.8%. This indicates that LLMs prefer answers provided by CoE even when confronted with numerous conflicting documents in the external knowledge base. The most advanced LLM like GPT-4o show particularly strong preference for CoE content, reaching an average ACC of 90.0%, due to their superior reasoning capabilities and ability to recognize the tight logical connections and question-relevant evidence within CoE structures.
Table 3: Effectiveness of different settings for Authority and CoE components (values reported in %).
@@ 6 Conclusion
In this paper, we present AuthChain, a novel single document poisoning attack method that leverages Chain-of-Evidence theory and authority effect to craft more convincing poisoned documents for RAG systems. Through extensive experiments on six popular LLMs, we demonstrate that AuthChain achieves significantly higher ASR while maintaining superior stealthiness compared to state-of-the-art baselines, despite injecting only one poisoned document per question. Our findings highlight the importance of developing more robust defense mechanisms for knowledge base security in RAG systems.
@@
Appendix A Example of CoE and CoE Features
Figure 3: Example of CoE and the CoE features.
Taking Figure 3 as an example, intent specifies “state location of business” as the user question goal, indicating the user wants to find the state where the business operates. Evidence nodes are the key entities extracted from user question, i.e., “drug stores”, “CEO”, and “Warren Bryant”. These nodes serve as bridges to connect the question with external knowledge about “Longs Drugs Store Corporation”. Evidence relations show how these entities are linked, with “have” connecting “drug stores” to “CEO”, and “is” linking “CEO” to “Warren Bryant”. The effectiveness of CoE stems from the synergistic interaction of these three features. The integration of all three features creates a comprehensive evidence chain that forms a complete knowledge structure tailored to the specific question.
@@ Appendix B Example of Poisoned Document Generated by AuthChain
Appendix C Details of Information Extraction Prompts

### v2

@@ 4.4 Baselines
  • PoisonedRAG (PRAG) Zou et al. (2024): Uses LLM to generate supporting content prefixed with original questions to ensure high retrieval probability.
@@ 4.5 Experiment Design and Metric
For the approach implementation, we leverage GPT-4 as the backbone for the Intent Agent, the CoE Judge Agent, the revise agent, and the Authority Content Agent, consistent with the original PoisonedRAG work where GPT-4 was used to generate poisoned data Zou et al. (2024). The temperature parameter is set to 0.1 for all agents to ensure the stability and consistency of the generated content.
To answer RQ1, we investigate single document poisoning attacks where each method (AuthChain and baselines) constructs and injects one poisoned document per targeted question. We evaluate their effectiveness in manipulating RAG systems’ outputs and analyze AuthChain’s performance against baselines, while also examining its internal components (CoE and Authority content) for comprehensive analysis.
@@
To answer RQ2, we select two representative RAG defense frameworks designed to counteract knowledge poisoning attacks: InstructRAG Wei et al. (2024) and AstuteRAG Wang et al. (2024b). InstructRAG enhances the robustness of RAG systems by explicitly guiding language models to learn a denoising process based on self-synthesized rationales. In this framework, the model is instructed to explain how the ground-truth answer is derived from the retrieved documents. These rationales can be leveraged as in-context demonstrations for explicit denoising or as supervised fine-tuning data, thereby improving the model’s ability to identify and resist poisoned or misleading knowledge in the retrieval set. AstuteRAG improves the robustness of RAG systems against imperfect or malicious retrieval by analyzing conflicts between the LLM’s internal knowledge and external sources. It adaptively extracts key information from internal knowledge, integrates it with retrieved content, and produces answers based on source reliability. This method has shown strong effectiveness in detecting and mitigating knowledge poisoning attacks. For fair comparison, we constrain all attacks (both AuthChain and baselines) to inject only a single poisoned document, and evaluate them under these two defense frameworks against the clean setting where no poisoned document is injected.
To answer RQ3, we conduct two experiments: (1) Authority Setting: Our goal is to examine whether authority-enhanced documents can influence LLM decisions even when they conflict with the LLM’s internal knowledge. To create an effective test environment, we need cases where LLMs have internal knowledge about the answers. We sample 600 QA pairs from HotpotQA and identify 118 questions that GPT-3.5 can correctly answer without external retrieval, indicating strong internal knowledge. These questions serve as ideal test cases since later GPT models also retain this internal knowledge. For these questions, we first create poisoned documents by modifying the correct answers in authentic documents to incorrect ones (Raw documents), then enhance these poisoned documents with authority signals using AuthChain. To simulate real-world scenarios where correct and incorrect information coexist, we gradually introduce authentic documents containing correct answers, creating mixed knowledge bases with correct document proportions (CDP) of 0.5, 0.67, and 0.75.
(2) CoE Setting: Our goal is to evaluate whether LLMs show stronger preference for CoE documents over conflicting information in the retrieved context. From the same 600 QA pairs, we identify 323 supporting documents that contain evidence for correct answers but lack structured evidence chains (Raw documents). We transform these into CoE-structured documents using AuthChain. To create challenging test scenarios, we introduce GPT-4 generated poisoned documents containing evidence for incorrect answers. We create mixed knowledge bases with poisoned document proportions (PDP) of 0.5, 0.67, and 0.75.
We evaluate AuthChain using four metrics: Attack Success Rate (ASR), Retrieval Success Rate (RSR), Perplexity (PPL), and Accuracy (ACC). For evaluating the poisoning effectiveness (RQ1), we measure ASR as the proportion of questions where the LLM’s output contains the answer from the poisoned target document. Following previous works Rizqullah et al. (2023); Huang et al. (2023), we determine the presence of target answers using substring matching. We also examine RSR, which represents the proportion of poisoned target documents successfully retrieved among the top-5 documents, and PPL (calculated using GPT-2 Radford et al. (2019)) to measure text fluency where higher values indicate less natural text. For defense evaluation (RQ2), we compare both ASR and ACC, where ACC reflects the proportion of questions where the LLM’s response contains the correct answer. In RQ3, we investigate whether authority-enhanced documents can overcome LLMs’ internal knowledge bias by comparing ASR across different CDP, and whether CoE-structured documents are more influential than raw documents when competing with conflicting information by comparing ACC across different PDP.
@@ 5.1 Answering RQ1
Dataset Metric Model PRAG HRAG CoE Authority AuthChain
HotpotQA ASR GPT-3.5 69.0 57.0 79.0 82.0 90.0
@@
RSR - 97.0 87.0 93.0 65.0 94.0
PPL - 73.8 398.0 28.5 49.5 26.4
We evaluate the effectiveness of AuthChain against six LLMs (GPT-3.5, GPT-4, GPT-4o, Llama3-8B, Llama3-70B, and DeepSeek-V3) across three widely-used datasets (HotpotQA, MS-MARCO, and NQ). Table 1 compares the Attack Success Rate (ASR) of AuthChain with both external baselines (PRAG and HRAG) and internal components (CoE and Authority content), along with key metrics including Retrieval Success Rate (RSR) and perplexity (PPL).
AuthChain achieves an average ASR of 87.0%, 81.5%, and 77.8% on HotpotQA, MS-MARCO, and NQ respectively, surpassing PRAG (21.7%-36.7%) and HRAG (31.0%-46.5%) by a significant margin. The results demonstrate the superior effectiveness of AuthChain on the single document poisoning scenario, which stems from our principle-guided content generation that better aligns with questions’ logical and semantic structure, further strengthened by authority expressions that enhance content credibility.
Specifically, while baselines resort to directly copying questions into poisoned documents, AuthChain dynamically generates content by incorporating key question elements through a principled three-stage design. The intent-based content generation successfully aligns with the question objective, while CoE content generation strengthens the logical alignment with question through constructed complete evidence chains. The effectiveness is further enhanced by our authority content generation which effectively leverages LLMs’ trust in authoritative sources. This structured approach generates more natural and authentic content while maintaining competitive retrievability, achieving only an average 2.3% decrease in RSR compared to PRAG and 5.0% higher than HRAG. The superior content quality is quantitatively validated by significantly lower perplexity scores (average PPL of 33.4 versus 69.9 for PRAG and 381.1 for HRAG).
Our component analysis reveals the effectiveness of both CoE and Authority content. Through an efficient generation process where judge agent and revise agent iterate only 1.3 times on average, our CoE content constructs complete evidence chains that preserve question structure, achieving 10.2%-18.8% higher ASR than baselines. This demonstrates its ability to efficiently compete with fragmented external knowledge while maintaining strong retrievability, with RSR only 4.0% lower than PRAG but 4.4% higher than HRAG. The Authority content shows 10.0%-18.6% ASR improvement through incorporating domain-specific authority signals, revealing LLMs’ susceptibility to authority bias across different architectures and scales. When combined, AuthChain surpasses CoE and authority content by 18.4% and 18.6% respectively, indicating strong complementarity: CoE content ensures competitive retrieval and dominance over external knowledge, while Authority content enhances trustworthiness to overcome LLMs’ internal knowledge.
Besides, to alleviate the computational costs of using GPT-4 based agents, we explore open-source LLMs as alternative agents in Appendix E. The results show these alternatives achieve comparable performance while being more cost-effective.
@@ 5.3 Answering RQ3
Table 2: Effectiveness of different settings for Authority and CoE content (values reported in %).
@@
Table 2 presents results from two distinct experiments examining the effectiveness of Authority and CoE content. The first experiment investigates the Attack Success Rate (ASR) with and without authority content as the Correct Document Proportion (CDP) increases. The second experiment evaluates the impact of CoE on LLM accuracy (ACC) under varying Poisoned Document Proportion (PDP).
For the authority setting, with raw poisoned documents (without authority content) and CDP set to 0.5, LLMs show considerable internal robustness, resulting in relatively low ASR (average 32.2%). This is particularly evident in advanced models like GPT-4o, where the ASR is only 17.9%, demonstrating strong internal knowledge resistance to poisoning. As CDP increases, the ASR of raw poisoned documents drops even further. In contrast, incorporating authority content significantly improves attack effectiveness, achieving an average ASR of 60.9% with CDP of 0.5, marking a 28.7% increase over raw poisoned documents. Moreover, as CDP rises to 0.75, documents with authority statements maintain better effectiveness, showing only a 15.6% ASR decrease compared to the 23.3% decrease in raw documents. This demonstrates that authority statements effectively overcome LLMs’ internal knowledge barriers and enhance the credibility of poisoned content.
Regarding the CoE setting, supporting documents structured with CoE help LLMs achieve an average ACC of 88.1% compared to 79.6% with raw supporting documents when PDP is 0.5. Even as PDP rises to 0.75, CoE supporting documents maintain an ACC of 83.2%, while raw documents drop to 72.8%. This indicates that LLMs prefer answers provided by CoE even when confronted with mutiple conflicting documents in the external knowledge base.
@@ 6 Conclusion
In this paper, we present AuthChain, a realistic knowledge poisoning attack that achieves successful attacks by poisoning only a single document while remaining effective for complex multi-hop questions. AuthChain addresses three key challenges to ensure the poisoned documents are reliably retrieved and trusted by the LLM, even against large knowledge bases and LLM’s own knowledge. Through extensive experiments on six popular LLMs, we demonstrate that AuthChain achieves significantly higher attack success rates while maintaining superior stealthiness against RAG defense mechanisms compared to state-of-the-art baselines.
@@
Ethical Statement
Our research on knowledge base poisoning attacks is conducted solely for defensive purposes and academic understanding. We have strictly followed responsible disclosure practices by promptly reporting our findings to relevant knowledge base platforms and RAG system providers. To prevent potential misuse, we only present necessary methodological insights and proof-of-concept examples in this paper, withholding specific attack implementations until appropriate defense mechanisms are developed. This work aims to enhance RAG system security rather than enable harmful exploits.
Appendix A Example of Question-Derived Features and LLM Preferred Knowledge
Figure 3: Question-derived features and examples of LLM preferred knowledge.
Taking Figure 3 as an example, intent specifies “state location of business” as the user question goal, indicating the user wants to find the state where the business operates. Evidence nodes are the key entities extracted from user question, i.e., “drug stores”, “CEO”, and “Warren Bryant”. These nodes serve as bridges to connect the question with external knowledge about “Longs Drugs Store Corporation”. Evidence relations show how these entities are linked, with “have” connecting “drug stores” to “CEO”, and “is” linking “CEO” to “Warren Bryant”. The integration of all question-derived features creates a comprehensive evidence chain that forms a complete knowledge structure tailored to the specific question.
@@ Appendix B Example of Poisoned Document Generated by AuthChain
Appendix C The Result for Answering RQ2
Dataset Model InstructRAG AstuteRAG
PRAG HRAG AuthChain Clean PRAG HRAG AuthChain Clean
[…]
HotpotQA GPT-3.5 49.0/42.0 40.0/46.0 36.0/60.0 76.0 59.0/33.0 61.0/25.0 46.0/52.0 78.0
GPT-4 56.0/38.0 52.0/39.0 47.0/52.0 79.0 79.0/10.0 74.0/13.0 63.0/28.0 78.0
GPT-4o 68.0/31.0 70.0/24.0 58.0/40.0 81.0 72.0/11.0 78.0/10.0 52.0/40.0 79.0
Llama3-8B 53.0/42.0 48.0/44.0 47.0/50.0 82.0 70.0/15.0 52.0/31.0 46.0/48.0 70.0
Llama3-70B 66.0/30.0 78.0/14.0 60.0/38.0 84.0 72.0/17.0 76.0/8.0 51.0/45.0 83.0
Deepseek-V3-0324 69.0/30.0 62.0/32.0 60.0/36.0 80.0 78.0/8.0 71.0/15.0 60.0/34.0 75.0
MS-MARCO GPT-3.5 51.0/40.0 47.0/43.0 45.0/46.0 78.0 73.0/14.0 81.0/12.0 56.0/30.0 82.0
GPT-4 65.0/30.0 59.0/39.0 56.0/43.0 86.0 88.0/5.0 85.0/8.0 57.0/34.0 89.0
GPT-4o 76.0/21.0 81.0/13.0 70.0/25.0 82.0 86.0/4.0 84.0/8.0 67.0/28.0 87.0
Llama3-8B 56.0/37.0 49.0/48.0 51.0/46.0 83.0 83.0/11.0 73.0/23.0 60.0/35.0 91.0
Llama3-70B 67.0/29.0 71.0/22.0 50.0/48.0 79.0 86.0/9.0 58.0/39.0 56.0/37.0 91.0
Deepseek-V3-0324 82.0/16.0 72.0/20.0 69.0/28.0 89.0 88.0/5.0 85.0/10.0 82.0/11.0 89.0
NQ GPT-3.5 49.0/42.0 45.0/50.0 41.0/56.0 66.0 62.0/21.0 75.0/9.0 44.0/47.0 71.0
GPT-4 64.0/29.0 54.0/40.0 51.0/44.0 75.0 81.0/5.0 83.0/7.0 69.0/23.0 80.0
GPT-4o 83.0/14.0 75.0/18.0 69.0/27.0 78.0 80.0/8.0 82.0/9.0 71.0/23.0 84.0
Llama3-8B 54.0/38.0 50.0/40.0 49.0/44.0 74.0 78.0/9.0 69.0/17.0 60.0/33.0 78.0
Llama3-70B 65.0/30.0 72.0/16.0 47.0/45.0 79.0 83.0/5.0 87.0/5.0 58.0/38.0 86.0
Deepseek-V3-0324 79.0/21.0 69.0/28.0 68.0/30.0 82.0 86.0/0.0 84.0/6.0 75.0/11.0 86.0
Appendix D Details of Information Extraction Prompts